According to the midooint methed, the price elastility of demand for apeles between point X and point Y is aprmatlywhich suopsts tha the demana ser .eekesis-.▼ between ponts x and Y.
Answer #1

From the graph:

P1: $1

P2: $2

Q1: 80 pounds

Q2: 70 Pounds

easlticity using midpoint formula:

e = %change in quantity/% change in price

%change in quantity: (Q2-Q1)/Q1+Q2/2

= -10/75

%change in price : (P2-P1)/P1+P2/2

= 1/1.5

hence e = -10*1.5/75*1 = -0.2

Hence elasticity of demand of apples is relatively inelastic. (−1 < e< 0)